# Adaptive Width Neural Networks

### v4

@@ 6 Conclusions
We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons. Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence. We showed stability of convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space. A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters. Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
Appendix A Background Notions
Since we were mostly inspired by the work of Nazaret and Blei [42], we complement the main text with an introduction to truncated distributions. In particular, by truncating a distribution at its quantile function evaluated at $k$ (see Appendix B for a visual explanation), the support of the distribution becomes finite and countable, hence we can compute expectations in a finite number of steps. A truncated distribution should ideally satisfy some requirements defined in Nazaret and Blei [42] and illustrated below.
Definition A.1.
A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
@@ Appendix A Background Notions
    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
@@
An important distinction with Nazaret and Blei [42] is that we do not need nor want condition 2 to be satisfied. As a matter of fact, we want to enforce a strict ordering of neurons where the first is always the most important one. Condition 2 is useful, as done in Nazaret and Blei [42], to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result. In the context of adaptive-width networks, however, this mechanism may leave a lot of neurons completely unutilized, letting the network grow indefinitely. That is why we require that each distribution $q$ is monotonically decreasing and all possible widths should have non-zero importance. We change the definition for the distributions we are interested in as follows:
@@
A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
@@
    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
@@
    $\forall q\in Q$, $q$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},q(x)>0$
@@
    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
@@ Appendix C Full ELBO Derivation
$\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right]=$ (11)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[\log p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)-\log q(\boldsymbol{\lambda},\boldsymbol{\theta})]=$ (12)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log(p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)p(\boldsymbol{\theta})p(\boldsymbol{\lambda}))-\log(q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda}))]$ (13)
@@
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)+\log p(\boldsymbol{\lambda})+\log p(\boldsymbol{\theta})-\log q(\boldsymbol{\lambda})-\log q(\boldsymbol{\theta}|\boldsymbol{\lambda})]=$ (14)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)]\approx$ (15)
$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)]$ (16)
where we use the first-order approximation in the last step. Note that when computing $\frac{p(\theta)}{q(\theta|\lambda)}$, the products from $D_{\ell+1}$ to infinity cancel out. Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)=\prod_{i}^{N}p(y_{i}|\boldsymbol{\nu},\boldsymbol{\rho},x_{i})$, and by transforming products into summations using the logarithm.
An important note about the first-order (Taylor) approximation $E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})q(\boldsymbol{\theta}|\boldsymbol{\lambda};\boldsymbol{\rho})}[f(\boldsymbol{\lambda},\boldsymbol{\theta})]\approx E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})}[f(\boldsymbol{\lambda},\boldsymbol{\rho})]\approx f(\boldsymbol{\nu},\boldsymbol{\rho})$ is that the function $f(\boldsymbol{\lambda},\boldsymbol{\theta})=\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it. This also happens in Nazaret and Blei [42], Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it. Importantly, the function is not differentiable at a countable set of points (one for each possible width), which has a measure null. Since the Lebesgue integral of the expectation is over reals, we should be able to remove this set from the integral while ensuring that the approximation still holds mathematically.
@@ Appendix D Lipschitz Continuity of the ELBO
Whenever the quantile of the finite distribution changes, we need to change the number of neurons of the neural network. Here, we want to show that the change in the ELBO after a change in the number of neurons is bounded. Therefore, we provide a theoretical result (Proposition D.1) that shows that the ELBO satisfies the Lipschitz continuity.
Proposition D.1 (AWNN ELBO Lipschitz continuity).
The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $\ell$, is Lipschitz continuous.
@@
We focus on the term involving $D_{\ell}$ of the ELBO, we write Equation 9 as
$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ Let’s first define
$\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
@@
$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (17)
$\leq~\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
$\leq\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
@@
with $M=\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$. We now look at the last term,
$f_{2}(\mathbf{D})=\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
This function is a neural network followed by the squared norm. Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere [59], therefore Lipschitz on both parameters and the input. When we change $D_{\ell}$ to $D_{\ell^{\prime}}$, we are adding network parameters, in particular, if we look at the two-layer network, we have
@@ Appendix E Full derivation of the rescaled weight initialization
For reasons that will become clear later, we may want to compute the gradient of the loss function with respect to the intermediate activations $\boldsymbol{\alpha}^{\ell}$ at a given layer, that is $\nabla\mathcal{L}(\boldsymbol{\alpha}^{\ell})=\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell})}{\partial\alpha_{1}^{\ell}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell}}^{\ell})}{\partial\alpha_{N^{\ell}}^{\ell}}\right)$. We focus on the $i$-th partial derivative $\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}$, where the only variable is $\alpha_{i}^{\ell}$. Then, we view the computation of the loss function starting from $\alpha_{i}^{\ell}$ as a composition of a function $\boldsymbol{\alpha}^{\ell+1}:\mathbb{R}\rightarrow\mathbb{R}^{N^{\ell+1}}=\left(\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})\dots,\alpha_{{N^{\ell+1}}}^{\ell+1}(\alpha_{i}^{\ell})\right)$ and another function (abusing the notation) $\mathcal{L}:\mathbb{R}^{N^{\ell+1}}\rightarrow\mathbb{R}$ that computes the loss value starting from $\boldsymbol{\alpha}^{\ell+1}$. By the chain rule:
@@
Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot and Bengio [20]
@@
Finally, because the mean of the independent variables involved is zero by assumption, it holds that $\mathrm{Var}[w^{\ell}_{ij}\alpha^{\ell-1}_{j}]=\mathrm{Var}[w^{\ell}_{ij}]\mathrm{Var}[\alpha^{\ell-1}_{j}]$. We can also abstract from the indexes, since the weight variables are i.i.d. and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^4] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot and Bengio [20], since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$. In Glorot and Bengio [20] our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
@@
From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^5] Note that what we impose is different from Glorot and Bengio [20], where the specific position $i$ is irrelevant. Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant. Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
Using Equation 22, and considering as in Glorot and Bengio [20] that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
@@
[…]
@@
From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in He et al. [25]: we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability. One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$. As a result, one can re-write Equation 26 as
@@
Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").

### v5

@@ 6 Conclusions
We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons. Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence. We showed stability of convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space. A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters. Finally, we have tested AWN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
7 Reproducibility Statement
The code provided in the supplementary material relies on libraries for automatic experimentation, which enforce reproducibility of the results. These libraries require that the user specifies every detail, from data splitting strategies to hyper-parameter selection and final evaluations, and subsequently the experiment is automatized and results are provided.
Appendix A Truncated Distributions’ Notions
Since we were mostly inspired by the work of Nazaret & Blei (2022), we complement the main text with an introduction to truncated distributions. In particular, by truncating a distribution at its quantile function evaluated at $k$ (see Appendix B for a visual explanation), the support of the distribution becomes finite and countable, hence we can compute expectations in a finite number of steps. A truncated distribution should ideally satisfy some requirements defined in Nazaret & Blei (2022) and illustrated below.
Definition A.1 (Nazaret & Blei (2022)).
A variational family $Q={q(x;\bm{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
@@ Appendix A Truncated Distributions’ Notions
    Each parameter in the set $\bm{\omega}$ is a continuous variable.
@@
An important distinction with Nazaret & Blei (2022) is that we do not need nor want condition 2 to be satisfied. As a matter of fact, we want to enforce a strict ordering of neurons where the first is always the most important one. Condition 2 is useful, as done in Nazaret & Blei (2022), to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result. In the context of adaptive-width networks, however, this mechanism may leave a lot of neurons completely unutilized, letting the network grow indefinitely. That is why we require that each distribution $f_{\ell}$ is monotonically decreasing and all possible widths should have non-zero importance. We change the definition for the distributions we are interested in as follows:
@@
A family of distributions $Q={f(x;\bm{\lambda})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
@@
    $\forall f\in Q$, $\operatorname{support}(f)$ is bounded
@@
    $\forall f\in Q$, $f$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},f(x)>0$
@@
    Each parameter in the set $\bm{\lambda}$ is a continuous variable.
@@ Appendix B How to Compute $D_{\ell}$ in Practice
Note that typically, newly introduced neurons are multiplied by a very small importance, and their random weights are sampled from a standard Gaussian. This makes their initial contribution very small, so it does not drastically alter the loss/gradients/convergence. From a backpropagation viewpoint, the reason why new neurons are introduced is that previous neurons need to become relatively more important. So, when the chosen importance distributions grow too slowly, as in the case of a Power law family of distributions, this can slow down convergence because it takes a long time to make the last neurons more important, and in addition a large tail or unimportant neurons is added if the chosen quantile is too high. The exponential distribution, instead, worked pretty well without having to tune it extensively.
@@ Appendix C Full ELBO Derivation
$\mathbb{E}_{q(\bm{\lambda},\bm{\theta})}\left[\log\frac{p(Y,\bm{\lambda},\bm{\theta}|X)}{q(\bm{\lambda},\bm{\theta})}\right]=$ (11)
$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta}|\bm{\lambda})}[\log p(Y,\bm{\lambda},\bm{\theta}|X)-\log q(\bm{\lambda},\bm{\theta})]=$ (12)
$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log(p(Y|\bm{\lambda},\bm{\theta},X)p(\bm{\theta})p(\bm{\lambda}))-\log(q(\bm{\lambda})q(\bm{\theta}|\bm{\lambda}))]$ (13)
@@
$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log p(Y|\bm{\lambda},\bm{\theta},X)+\log p(\bm{\lambda})+\log p(\bm{\theta})-\log q(\bm{\lambda})-\log q(\bm{\theta}|\bm{\lambda})]=$ (14)
$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log\frac{p(\bm{\lambda})}{q(\bm{\lambda})}+\log\frac{p(\bm{\theta})}{q(\bm{\theta}|\bm{\lambda})}+\log p(Y|\bm{\lambda},\bm{\theta},X)]\approx$ (15)
$\log\frac{p(\bm{\nu})}{q(\bm{\nu})}+\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}+\log p(Y|\bm{\nu},\bm{\rho},X)]$ (16)
where we use the first-order approximation in the last step. Note that when computing $\frac{p(\theta)}{q(\theta|\lambda)}$, the products from $D_{\ell+1}$ to infinity cancel out. Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\bm{\nu},\bm{\rho},X)=\prod_{i}^{N}p(y_{i}|\bm{\nu},\bm{\rho},x_{i})$, and by transforming products into summations using the logarithm.
An important note about the first-order (Taylor) approximation $E_{q(\bm{\lambda};\bm{\nu})q(\bm{\theta}|\bm{\lambda};\bm{\rho})}[f(\bm{\lambda},\bm{\theta})]\approx E_{q(\bm{\lambda};\bm{\nu})}[f(\bm{\lambda},\bm{\rho})]\approx f(\bm{\nu},\bm{\rho})$ is that the function $f(\bm{\lambda},\bm{\theta})=\log\frac{p(\bm{\lambda})}{q(\bm{\lambda})}+\log\frac{p(\bm{\theta})}{q(\bm{\theta}|\bm{\lambda})}+\log p(Y|\bm{\lambda},\bm{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it. This also happens in Nazaret & Blei (2022), Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it. Importantly, the function is not differentiable at a countable set of points (one for each possible width), which has a measure null. Since the Lebesgue integral of the expectation is over reals, we should be able to remove this set from the integral while ensuring that the approximation still holds mathematically.
@@ Appendix D Lipschitz Continuity of the ELBO
Whenever the quantile of the finite distribution changes, we need to change the number of neurons of the neural network. Here, we want to show that the change in the ELBO after a change in the number of neurons is bounded. Therefore, we provide a theoretical result (proposition˜D.1) that shows that the ELBO satisfies the Lipschitz continuity.
Proposition D.1 (AWN ELBO Lipschitz continuity).
The ELBO loss of eq.˜9, with respect to the change in the depth $D_{\ell}$ for the layer $\ell$, is Lipschitz continuous.
@@
We focus on the term involving $D_{\ell}$ of the ELBO, we write eq.˜9 as
$\log\frac{p(\bm{\nu})}{q(\bm{\nu})}+\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}+\log p(Y|\bm{\nu},\bm{\rho},X)$
where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ . Let’s first define
$\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\bm{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
@@
$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$ (17)
$\leq \sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$ (18)
$\leq\max_{n}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
@@
with $M=\max_{n}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$. We now look at the last term,
$f_{2}(\mathbf{D})=\log p(Y|\bm{\nu},\bm{\rho},X)$
This function is a neural network followed by the squared norm. Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere (Virmaux & Scaman, 2018), therefore Lipschitz on both parameters and the input. When we change $D_{\ell}$ to $D_{\ell^{\prime}}$, we are adding network parameters, in particular, if we look at the two-layer network, we have
@@ Appendix E Full derivation of the rescaled weight initialization
For reasons that will become clear later, we may want to compute the gradient of the loss function with respect to the intermediate activations $\bm{\alpha}^{\ell}$ at a given layer, that is $\nabla\mathcal{L}(\bm{\alpha}^{\ell})=\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell})}{\partial\alpha_{1}^{\ell}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell}}^{\ell})}{\partial\alpha_{N^{\ell}}^{\ell}}\right)$. We focus on the $i$-th partial derivative $\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}$, where the only variable is $\alpha_{i}^{\ell}$. Then, we view the computation of the loss function starting from $\alpha_{i}^{\ell}$ as a composition of a function $\bm{\alpha}^{\ell+1}:\mathbb{R}\rightarrow\mathbb{R}^{N^{\ell+1}}=\left(\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})\dots,\alpha_{{N^{\ell+1}}}^{\ell+1}(\alpha_{i}^{\ell})\right)$ and another function (abusing the notation) $\mathcal{L}:\mathbb{R}^{N^{\ell+1}}\rightarrow\mathbb{R}$ that computes the loss value starting from $\bm{\alpha}^{\ell+1}$. By the chain rule:
@@
Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot & Bengio (2010)
@@
Finally, because the mean of the independent variables involved is zero by assumption, it holds that $\mathrm{Var}[w^{\ell}_{ij}\alpha^{\ell-1}_{j}]=\mathrm{Var}[w^{\ell}_{ij}]\mathrm{Var}[\alpha^{\ell-1}_{j}]$. We can also abstract from the indexes, since the weight variables are i.i.d. and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^3] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot & Bengio (2010), since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$. In Glorot & Bengio (2010) our $D_{\ell-1}$ is denoted as $n_{i^{\prime}}$” (see Equation 5).
@@
From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^4] Note that what we impose is different from Glorot & Bengio (2010), where the specific position $i$ is irrelevant. Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant. Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
Using Equation 22, and considering as in Glorot & Bengio (2010) that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
@@
$\frac{1}{2}\int_{-\infty}^{{+\infty}}x^{2}p(x)dx=\frac{1}{2}\int_{-\infty}^{{+\infty}}(x-\mathbb{E}[x])^{2}p(x)dx=\frac{1}{2}\mathrm{Var}[x].$ (44)
@@
From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in He et al. (2015): we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability. One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$. As a result, one can re-write Equation 26 as
@@
Below, we provide a comparison of convergence between an AWN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").